OpenStudy (anonymous):

A basket contains the following pieces of fruit 3 apples 2 oranges 2 bananas 2 pears and 5 peaches. Jamson picks a fruit at random and does not replace it. Then Brittany picks a fruit at random. What is the probability that Jameson gets a banana and Brittany gets a pear?

OpenStudy (mathstudent55):

Find the probability of each drawing, then multiply them together.

OpenStudy (mathstudent55):

For the first drawing, you want a banana. How many bananas are there? How many pieces of fruit are there in total?

OpenStudy (anonymous):

14 pieces of fruit .. total

OpenStudy (mathstudent55):

Right. How many bananas?

OpenStudy (anonymous):

2 bananas

OpenStudy (mathstudent55):

Right. We can now do the probability of the first drawing. Probability of picking a banana in the first drawing is P(banana) = 2/14 = 1/7

OpenStudy (mathstudent55):

Ok?

OpenStudy (anonymous):

ok thank you

OpenStudy (mathstudent55):

We're not done yet.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

The second pick is a pear. Now we need to see what is the probability of picking a pear in the second drawing.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

There were 14 pieces of fruit at the beginning. A banana has been taken and has not been replaced. How many pieces of fruit are there now before the second drawing?

OpenStudy (anonymous):

13

OpenStudy (mathstudent55):

Good. There are still 2pears, so the probability of picking a pear in the second drawing is: P(pear) = 2/13

OpenStudy (anonymous):

right ok

OpenStudy (mathstudent55):

The probability of picking a banana followed by a pear is the product of the two probabilities: P(banana followed by pear) = P(banana) * P(pear) = 1/7 * 2/13 = 2/91

OpenStudy (mathstudent55):

The final answer is 2/91

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

You're welcome.